# Beyond the random phase approximation for calculating Curie temperatures in ferromagnets: application to Fe, Ni, Co and monolayer CrI3

### v1

@@ Abstract
The magnetic properties of solids are typically analyzed in terms of Heisenberg models where the electronic structure is approximated by interacting localized spins. However, even in such models the evaluation of thermodynamic properties constitutes a major challenge and is usually handled by a mean field decoupling scheme. The random phase approximation (RPA) comprises a common approach and is often applied to evaluate critical temperatures although it is well known that the method is only accurate well below the critical temperature. In the present work we compare the performance of the RPA with a different decoupling scheme proposed by Callen as well as the mean field decoupling of interacting Holstein-Primakoff (HP) magnons. We consider three-dimensional (3D) as well as two-dimensional (2D) model systems where the Curie temperature is governed by anisotropy. In 3D, the Callen method is the most accurate in the classical limit, and we show that the Callen decoupling produces the best agreement with experiments for bcc Fe, fcc Ni and fcc Co with exchange interactions obtained from first principles. In contrast, for low spin systems where a quantum mechanical treatment in pertinent, the HP and RPA methods appear are superior to the Callen decoupling. In 2D systems with magnetic order driven by single-ion anisotropy, it is shown that HP fails rather dramatically and both RPA and Callen approaches severely overestimates Curie temperatures. The most accurate approach is then constructed by combining RPA with the Callen decoupling of single-ion anisotropy, which yields the correct lack of order for $S=1/2$. We exemplify this by the case of monolayer CrI3 using exchange constant extracted from experiments.
@@ I Introduction
The phase transition associated with loss of magnetic order constitutes a defining property of any magnet and the critical temperature where it happens is thus a crucial descriptor for magnetic materials. Besides being of fundamental interest, the critical temperature typically determines whether or not a given material may be of technological relevance. For example, a prerequisite for developing new materials for either spintronics, sensors or generators is a critical temperature that lies well above room temperature. Nevertheless, the theoretical prediction of critical temperature still poses a major challenge and the reliability of calculations is often difficult to validate. In particular, first principles simulations are typically not able to include finite temperature effects and the analysis of thermodynamic properties are commonly carried out by mapping the electronic structure to a Heisenberg model [1, 2, 3, 4]. The problem is then reduced to solving a Hamiltonian of localized spins coupled through exchange interactions. For the case of $S=1/2$ the one-dimensional Heisenberg model can be diagonalized by the Bethe ansatz [5, 6] and for higher values of the spin, it may be treated accurately by the density matrix renormalization group [7]. In two or three dimensions it is more difficult to apply accurate numerical techniques to the Heisenberg model. It is straightforward to perform classical Monte Carlo simulations of the model and this approach seems to be gaining popularity in the literature [8, 9, 10, 11, 12]. It is, however, far from clear that the Curie temperatures obtained from the classical model comprise a reliable estimate. On the other hand, quantum Monte Carlo simulations [13] are computationally demanding and it is not obvious that the method is applicable to complex Heisenberg models with long-range exchange interactions obtained from first principles simulations using present day computational resources. Finally, there are several approximate mean field schemes that allow one to calculate thermodynamic properties of the Heisenberg model, but the accuracy of these are often difficult to assess. Such an assessment will be the main focus of the present work and we will show that commonly applied approximations are rather inaccurate and may lead to wrong validation of magnetic models due to spurious agreement of calculated Curie temperatures with experiments.
@@
A prominent difference between 2D and three-dimensional (3D) magnetism is that magnetic order in 2D cannot be driven by spontaneous symmetry breaking [33], but requires explicit symmetry breaking by magnetic anisotropy. The single-ion anisotropy energy is the simplest way to include magnetic anisotropy in the Heisenberg model [9], but the mean-field decoupling discussed in either of the methods mentioned above entails an ambiguity related to the ordering of operators prior to decoupling. Due to the increasing interest in 2D magnetism it is pertinent to resolve this ambiguity, which can have a strong influence on the model predictions. For example, single-ion anisotropy in 2D $S=1/2$ systems cannot induce magnetic order at finite temperatures and whether or not this result is obtained in either RPA or the HP method depends on an arbitrary choice of operator ordering prior to the mean field decoupling.
In the present work we perform a thorough comparison of the Callen method and compare with results obtained from RPA and HP decoupling as well as Weiss mean-field theory and classical Monte Carlo simulations. In particular we benchmark the methods in the classical limit where the exact result can be obtained numerically from Monte Carlo (MC) simulations. The Callen and HP methods are found to yield excellent agreement with MC, whereas RPA underestimates the critical temperatures. The approach is applied to the cases of fcc Ni, bcc Fe and fcc Co with exchange parameters obtained from DFT and it is shown that the Callen approach yields the best agreement with experiments for Ni and Co whereas HP seems to be superior for bcc Fe. We then analyze 2D magnetic order driven by magnetic anisotropy and discuss the subtleties related to decoupling single-ion anisotropy term. Finally, we discuss the case of monolayer CrI3 where we fix the exchange constant from the measured optical magnon energies. This value is in agreement with DFT calculations using the LDA functional and we argue that previous claims of accurate exchange constant from DFT+U (a factor two larger than LDA results) are due to a severe underestimation of the critical temperature from classical MC simulations.
@@ II Theory
where $\mathbf{S}_{i}$ is the spin operator of site $i$ positioned at the lattice vector $\mathbf{R}_{i}$ and summation over both $i$ and $j$ indices is implied. The inclusion of anisotropy terms will be considered later in section II.5. For ferromagnets where the spins are aligned along some quantization axis in the ground state—here chosen as the $z$-axis—it is furthermore convenient to introduce the spin raising and lowering operators
@@ II.2 Holstein-Primakoff approach
In the Holstein-Primakoff (HP) method [14], the spin operators are expressed in terms of bosonic operators satisfying $[a_{i},a^{\dagger}_{j}]=\delta_{ij}$. It may be shown that the commutator relations for the spin operators are maintained by the relations
$S_{i}^{+}$ $=\sqrt{2S-a_{i}^{\dagger}a_{i}}\,a_{i}$ (9a)
$S_{i}^{-}$ $=a_{i}^{\dagger}\sqrt{2S-a_{i}^{\dagger}a_{i}}$ (9b)
@@
where $S$ again is the maximum eigenvalue of $S^{z}$. Expanding the square roots in the Hamiltonian (3) and retaining only terms up to quadratic order, the problem becomes exactly solvable [34], generating a series of eigenstates that are interpreted as magnons. Higher order terms represent magnon-magnon interactions, which become important at elevated temperatures. Thus, for any accurate treatment of thermal properties these need to included and one is left with an intractable interacting bosonic Hamiltonian. Nevertheless, if the expansion is truncated at fourth order in bosonic operators, the interaction terms can be handled in a Hartree-Fock type of mean-field decoupling. This is most conveniently done in Fourier space. Defining
@@ II.3 Random phase approximation
The second term of Eq. (17a) can be expressed in terms of higher order Green’s functions that need to be approximated in order to obtain a closed expression for $G^{+-}_{ij}$. In the RPA, it is assumed that fluctuations in $S_{i}^{z}$ are small and it is replaced by its average value $\langle S^{z}\rangle$ (which is independent of the site index for ferromagnets with equivalent sites). This is expected to be true at low temperatures, but is probably not well justified when approaching the critical temperature. The central approximation is thus
@@ II.4 Callen decoupling
In essence the RPA neglects replaces $S^{z}$ with $\langle S^{z}\rangle$ and thus neglects fluctuations in $S^{z}$. While this is expected to be a good approximation at low temperatures it is likely to break down when the critical temperature is approached. To improve upon this, an alternative decoupling scheme was proposed by Callen [19]. The idea is based on the fact that different representations of the $S^{z}$ operator will lead to different mean field approximations for the Green’s function. Specifically, we may write $S^{z}$ in three distinct ways:
@@
Comparing with the RPA dispersion (21), we note the appearance of additional terms arising from the second decoupling term in Eq. (29). The Bose factor in this term introduces an additional temperature renormalization to the magnon energies similar to the HP dispersion (12). In fact, the only difference between Eqs. (12) and (31) is the prefactor $\langle S^{z}\rangle^{2}/S^{2}$, which is absent in HP dispersion. This factor, however, ensures that the additional renormalization of the RPA dispersion is decreased in weight as the Curie temperature is approached from below—in contrast to the HP approach where the renormalization leads to negative magnon energies at a finite value of $\langle S^{z}\rangle$. Hence, in a certain sense the Callen decoupling scheme can be viewed as an intermediary between the RPA and HP methods.
@@ II.5 Magnetic anisotropy
In the case of RPA, one faced with the question of how to decouple the single-ion anisotropy
@@ III Results
We will start by benchmarking the different decoupling schemes for isotropic 3D systems in the classical limit. For simplicity we begin with a cubic lattice with nearest neighbor interactions, where the results may be compared to classical Monte Carlo simulations. The methods are then applied to bcc Fe, fcc Ni and fcc Co with exchange constants obtained from first principles and the calculated critical temperatures are compared to experimental values. We then compare the methods for 2D ferromagnets, where anisotropy is decisive for obtaining finite Curie temperatures. We start by considering model Bravais lattices with nearest neighbor interactions and compare the classical limit to Monte Carlo simulations. Finally, we consider the case of CrI3 where we obtain the Curie temperature based on Heisenberg parameters extracted from experimental magnon energies. It will be argued that classical Monte Carlo simulations lead to severe underestimation of the critical temperature and accurate critical temperatures obtained from such approaches in the literature seem to rely on significant overestimation of calculated exchange constants.
@@ III.1 Cubic lattice with nearest neighbor isotropic interactions
Figure 1: The variation of critical temperature with spin for the cubic lattice with nearest neighbor isotropic exchange. The horizontal lines are the classical limits of each of the methods.
In Fig. 1 we compare the Curie temperature obtained with MFA, RPA, HP, CD as well as classical Monte Carlo (MC) simulations for the simple cubic lattice with nearest neighbor interactions. The results are shown as a function of $S$ and the Curie temperature is shown in units of $JS^{2}/k_{\mathrm{B}}$ such that the classical result is independent of $S$. We note that the spin dependence of the MFA and RPA results are given exactly by $S(S+1)$ and a similar dependence is roughly observed for HP and CD. At low values of $S$, RPA and HP is in agreement but start deviating at higher spin where CD and RPA starts converging towards similar values. In the classical limit ($S\rightarrow\infty$) the CD approach comes very close to the exact result (here represented by MC) whereas the HP overestimates the Curie temperature by 5%, RPA underestimates it by 10% and MFA overestimates it by 40%.
@@
In Ref. [35], QMC calculations for $S$=1/2 were performed and they observed that RPA has better agreement with the exact results than CD. In addition, Callen argued in his original paper that the method is probably not accurate for low spin. However, the CD method appears to be the most accurate for the classical limit, where it represents a significant simplification compared to performing classical MC that needs to be carefully converged and may comprise a highly demanding step in terms of computational load. The classical limit is often most appropriate for metals (to the extent that these may be modelled by a Heisenberg model), but not for insulators. In this regard, we note that the ratio between the MFA and RPA Curie temperatures calculated on the basis of the quantum mechanical and classical Heisenberg models respectively is given by $T_{\mathrm{c}}^{\mathrm{QM}}/T_{\mathrm{c}}^{\mathrm{Cl}}=(S+1)/S$. Commonly applied classical approaches (like MC [9]) are thus expected to underestimate the critical temperature severely - by a factor of three for $S=1/2$ and a factor $5/3$ for $S=3/2$.
@@ III.2.1 MFT calculations
where $\chi^{+-}_{\mathrm{KS}}$ is the static transverse Kohn-Sham susceptibility and $B_{i}^{\mathrm{xc}}$ is the exchange-correlation magnetic field truncated in a volume that defines the magnetic site. That is,
@@
In Fig. 2 we show the magnon dispersion relations of Fe, Ni and Co obtained from the bare as well as the renormalized expressions. The calculations were carried out with the electronic structure package GPAW [38] using the LDA functional. The q-point grids used for the MFT calculations were $\Gamma$-centered and uniform with $N_{q}^{1/3}=48$ for Fe and Ni and $N_{q}^{1/3}=36$ for Co. All calculations were done with a plane-wave cutoff of 1000 eV and 9 empty-shell bands. We used site volumes $\Omega_{i}$ centered on atoms with a radius of half the nearest neighbor distance. This yields site spins of 1.117, 0.3249, 0.8345 (in units of $\hbar$) and average exchange splittings $\delta^{\mathrm{xc}}$ of 2.3767, 0.79239 and 1.8769 eV for Fe, Ni and Co respectively. The site moments may be compared with the total spin per unit cell, which was 1.10, 0.314 and 0.815 respectively. The difference between the local and unit cell spin may be regarded as an inherent uncertainty in the method, which crucially relies in mapping an itinerant system to a localized spin model. The calculated magnon dispersion exhibit a significant effect of the renormalization in Ni and Co due to the rather small exchange splitting in these compounds, whereas the renormalization is less pronounced in Fe. The renormalized MFT dispersions are in excellent agreement with magnons energies obtained from time-dependent density functional theory (TDDFT) [39], whereas the bare MFT expressions underestimate the TDDFT results. The actual magnetic response of these systems is prone to rather strong Landau damping—in particular when approaching the Brillouin zone boundary—and the definition of a magnon energy from the TDDFT spectral function becomes ambiguous in some parts of the Brillouin zone [39]. The MFT approach will, however, by construction yield a magnon energy at any $\mathbf{q}$ and the noisy features of the magnon dispersions of Fig. 2 are related to nontrivial (not just Lorentzian broadening) interactions with the Stoner continuum. Nevertheless, we may use the MFT magnon energies to evaluate critical temperatures based on the different decoupling schemes discussed above, keeping in mind that the Heisenberg modelling of such systems is rather severely limited by the itinerant nature of the compounds.

### v2

@@ Abstract
The magnetic properties of solids are typically analyzed in terms of Heisenberg models where the electronic structure is approximated by interacting localized spins. However, even in such models the evaluation of thermodynamic properties constitutes a major challenge and is usually handled by a mean field decoupling scheme. The random phase approximation (RPA) comprises a common approach and is often applied to evaluate critical temperatures although it is well known that the method is only accurate well below the critical temperature. In the present work we compare the performance of the RPA with a different decoupling scheme proposed by Callen as well as the mean field decoupling of interacting Holstein-Primakoff (HP) magnons. We consider three-dimensional (3D) as well as two-dimensional (2D) model systems where the Curie temperature is governed by anisotropy. In 3D, the Callen method is the most accurate in the classical limit, and we show that the Callen decoupling produces the best agreement with experiments for bcc Fe, fcc Ni and fcc Co with exchange interactions obtained from first principles. In contrast, for low spin systems where a quantum mechanical treatment is pertinent, the HP and RPA methods are superior to the Callen decoupling. In 2D systems with magnetic order driven by single-ion anisotropy, it is shown that HP fails rather dramatically and both RPA and Callen approaches severely overestimates Curie temperatures. The most accurate approach is then constructed by combining RPA with the Callen decoupling of single-ion anisotropy, which yields the correct lack of order for $S=1/2$. We exemplify this by the case of monolayer CrI3 using exchange constant extracted from experiments.
@@ I Introduction
The phase transition associated with loss of magnetic order constitutes a defining property of any magnet and the critical temperature where it happens is thus a crucial descriptor for magnetic materials. Besides being of fundamental interest, the critical temperature typically determines whether or not a given material may be of technological relevance. For example, a prerequisite for developing new materials for either spintronics, sensors or generators is a critical temperature that lies well above room temperature. Nevertheless, the theoretical prediction of critical temperature still poses a major challenge and the reliability of calculations is often difficult to validate. In particular, first principles simulations are typically not able to include finite temperature effects and the analysis of thermodynamic properties are commonly carried out by mapping the electronic structure to a Heisenberg model [1, 2, 3, 4]. The problem is then reduced to solving a Hamiltonian of localized spins coupled through exchange interactions. For the case of $S=1/2$ the one-dimensional Heisenberg model can be diagonalized by the Bethe ansatz [5, 6] and for higher values of the spin, it may be treated accurately by the density matrix renormalization group [7]. In two or three dimensions it is more difficult to apply accurate numerical techniques to the Heisenberg model. It is straightforward to perform classical Monte Carlo simulations of the model and this approach seems to be gaining popularity in the literature [8, 9, 10, 11, 12]. It is, however, far from clear that the Curie temperatures obtained from the classical model comprise reliable estimates. On the other hand, quantum Monte Carlo simulations [13] are computationally demanding and it is not obvious that the method is applicable to complex Heisenberg models with long-range exchange interactions obtained from first principles simulations using present day computational resources. Finally, there are several approximate mean field schemes that allow one to calculate thermodynamic properties of the Heisenberg model, but the accuracy of these are often difficult to assess. Such an assessment will be the main focus of the present work and we will show that commonly applied approximations are rather inaccurate and may lead to wrong validation of magnetic models due to spurious agreement of calculated Curie temperatures with experiments.
@@
A prominent difference between 2D and three-dimensional (3D) magnetism is that magnetic order in 2D cannot be driven by spontaneous symmetry breaking [33], but requires explicit symmetry breaking by magnetic anisotropy. The single-ion anisotropy energy is the simplest way to include magnetic anisotropy in the Heisenberg model [9], but the mean-field decoupling discussed in either of the methods mentioned above entails an ambiguity related to the ordering of operators prior to decoupling. Due to the increasing interest in 2D magnetism it is pertinent to resolve this ambiguity, which can have a strong influence on the model predictions. For example, single-ion anisotropy in 2D $S=1/2$ systems cannot induce magnetic order at finite temperatures and whether or not this result is obtained in RPA depends on an arbitrary choice of operator ordering prior to the mean field decoupling.
In the present work we perform a thorough comparison of the Callen method and compare with results obtained from RPA and HP decoupling as well as Weiss mean-field theory and classical Monte Carlo simulations. In particular, we benchmark the methods in the classical limit where the exact result can be obtained numerically from Monte Carlo (MC) simulations. The Callen and HP methods are found to yield excellent agreement with MC, whereas RPA underestimates the critical temperatures. The approach is applied to the cases of fcc Ni, bcc Fe and fcc Co with exchange parameters obtained from DFT and it is shown that the Callen approach yields the best agreement with experiments for Ni and Co whereas HP seems to be superior for bcc Fe. We then analyze 2D magnetic order driven by magnetic anisotropy and discuss the subtleties related to mean field decoupling of the single-ion anisotropy term. Finally, we discuss the case of monolayer CrI3 where we fix the exchange constant from the measured optical magnon energies. The value disagrees with the previous claims of accurate exchange constant from DFT+U (a factor two larger than LDA results) and we argue that this is due to a severe underestimation of the critical temperature from classical MC simulations.
@@ II Theory
where $\mathbf{S}_{i}$ is the spin operator of site $i$ positioned at the lattice vector $\mathbf{R}_{i}$ and summation over both $i$ and $j$ indices is implied. The inclusion of anisotropy terms will be considered in section II.5. For ferromagnets where the spins are aligned along some quantization axis in the ground state—here chosen as the $z$-axis—it is furthermore convenient to introduce the spin raising and lowering operators
@@ II.2 Holstein-Primakoff approach
In the Holstein-Primakoff (HP) method [14], the spin operators are expressed in terms of bosonic operators satisfying $[a_{i},a^{\dagger}_{j}]=\delta_{ij}$. It may be shown that the commutator relations for the spin operators are conserved by the relations
$S_{i}^{+}$ $=\sqrt{2S-a_{i}^{\dagger}a_{i}}\,a_{i},$ (9a)
$S_{i}^{-}$ $=a_{i}^{\dagger}\sqrt{2S-a_{i}^{\dagger}a_{i}},$ (9b)
@@
where $S$ again denotes the maximum eigenvalue of $S^{z}$. Expanding the square roots in the Hamiltonian (3) and retaining only terms up to quadratic order, the problem becomes exactly solvable [34], generating a series of eigenstates that are interpreted as magnons. Higher order terms represent magnon-magnon interactions, which become important at elevated temperatures. Thus, for any accurate treatment of thermal properties these need to included and one is left with an intractable interacting bosonic Hamiltonian. Nevertheless, if the expansion is truncated at fourth order in bosonic operators, the interaction terms can be handled in a Hartree-Fock type of mean-field decoupling. This is most conveniently done in Fourier space. Defining
@@ II.3 Random phase approximation
The second term of Eq. (17a) can be expressed in terms of higher order Green’s functions that need to be approximated in order to obtain a closed expression for $G^{+-}_{ij}$. In the RPA, it is assumed that fluctuations in $S_{i}^{z}$ are small and it is replaced by its average value $\langle S^{z}\rangle$ (which is independent of the site index for ferromagnets with equivalent sites). The central approximation is thus
@@ II.4 Callen decoupling
In essence the RPA replaces $S^{z}$ with $\langle S^{z}\rangle$ and thus neglects fluctuations in $S^{z}$. While this is expected to be a good approximation at low temperatures it is likely to break down when the critical temperature is approached. To improve upon this, an alternative decoupling scheme was proposed by Callen [19]. The idea is based on the fact that different representations of the $S^{z}$ operator will lead to different mean field approximations for the Green’s function. Specifically, we may write $S^{z}$ in three distinct ways:
@@
Comparing with the RPA dispersion (21), we note the appearance of additional terms arising from the second decoupling term in Eq. (29). The Bose factor in this term introduces an additional temperature renormalization to the magnon energies similar to the HP dispersion (12). In fact, the only difference between Eqs. (12) and (31) is the prefactor $\langle S^{z}\rangle^{2}/S^{2}$, which is absent in HP dispersion. This factor, however, ensures that the additional renormalization of the RPA dispersion is decreased in weight as the Curie temperature is approached from below—in contrast to the HP approach where the renormalization leads to negative magnon energies at a finite value of $\langle S^{z}\rangle$. Hence, in a certain sense the Callen decoupling scheme can be viewed as intermediate between the RPA and HP methods.
@@ II.5 Magnetic anisotropy
In the case of RPA, one is faced with the question of how to decouple the single-ion anisotropy
@@ III Results
We will start by benchmarking the different decoupling schemes for isotropic 3D systems in the classical limit. For simplicity we begin with a cubic lattice with nearest neighbor interactions, where the results may be compared to classical MC simulations. The methods are then applied to bcc Fe, fcc Ni and fcc Co with exchange constants obtained from first principles and the calculated critical temperatures are compared to experimental values. We then compare the methods for 2D ferromagnets, where anisotropy is decisive for obtaining finite Curie temperatures. We start by considering model Bravais lattices with nearest neighbor interactions and compare the classical limit to Monte Carlo simulations. Finally, we consider the case of CrI3 where we obtain the Curie temperature based on Heisenberg parameters extracted from experimental magnon energies. It will be argued that classical Monte Carlo simulations lead to severe underestimation of the critical temperature and accurate critical temperatures obtained from such approaches in the literature seem to rely on significant overestimation of calculated exchange constants.
@@ III.1 Cubic lattice with nearest neighbor isotropic interactions
Figure 1: The variation of critical temperature with spin for the cubic lattice with nearest neighbor isotropic exchange. The horizontal lines are the classical limits for each of the methods.
In Fig. 1 we compare the Curie temperature obtained with MFA, RPA, HP, CD as well as classical Monte Carlo (MC) simulations for the simple cubic lattice with nearest neighbor interactions. All MC simulations are done by initializing a ferromagnetic configuration in a $20\times 20\times 20$ super cell. The critical temperature was obtained from the heat capacity which exhibits a peak at $T_{\mathrm{C}}$ and we checked that the results were converged with respect to super cell size and initial configuration of magnetic moments. The results are shown as a function of $S$ and the Curie temperature is shown in units of $JS^{2}/k_{\mathrm{B}}$ such that the classical result is independent of $S$. We note that the spin dependence of the MFA and RPA results are given exactly by $S(S+1)$ and a similar dependence is roughly observed for HP and CD. At low values of $S$, RPA and HP is in agreement but start deviating at higher spin where CD and RPA starts converging towards similar values. In the classical limit ($S\rightarrow\infty$) the CD approach comes very close to the exact result (here represented by MC) whereas the HP overestimates the Curie temperature by 5%, RPA underestimates it by 10% and MFA overestimates it by 40%.
@@
In Ref. [35], QMC calculations for $S$=1/2 revealed that RPA has better agreement with the exact results than CD. In addition, Callen argued in his original paper that the method is probably not accurate for low spin. However, the CD method appears to be the most accurate for the classical limit, where it represents a significant simplification compared to performing classical MC that needs to be carefully converged and may comprise a highly demanding step in terms of computational load. The classical limit is often most appropriate for metals (to the extent that these may be modelled by a Heisenberg model), but not for insulators. In this regard, we note that the ratio between the MFA and RPA Curie temperatures calculated on the basis of the quantum mechanical and classical Heisenberg models respectively is given by $T_{\mathrm{c}}^{\mathrm{QM}}/T_{\mathrm{c}}^{\mathrm{Cl}}=(S+1)/S$. Commonly applied classical approaches (like MC [9]) are thus expected to underestimate the critical temperature severely - by a factor of three for $S=1/2$ and a factor $5/3$ for $S=3/2$.
@@ III.2.1 MFT calculations
where $\chi^{+-}_{\mathrm{KS}}$ is the static transverse Kohn-Sham susceptibility and $B_{i}^{\mathrm{xc}}$ is the exchange-correlation magnetic field truncated in a volume that defines the magnetic site $i$. That is,
@@
In Fig. 2 we show the magnon dispersion relations of Fe, Ni and Co obtained from the bare as well as the renormalized expressions. The calculations were carried out with the electronic structure package GPAW [38] using the LDA functional. The q-point grids used for the MFT calculations were $\Gamma$-centered and uniform with $N_{q}^{1/3}=48$ for Fe and Ni and $N_{q}^{1/3}=36$ for Co. All calculations were done with a plane-wave cutoff of 1000 eV and 9 empty-shell bands. We used site volumes $\Omega_{i}$ centered on atoms with a radius of half the nearest neighbor distance. This yields site spins of 1.117, 0.3249, 0.8345 (in units of $\hbar$) and average exchange splittings $\delta^{\mathrm{xc}}$ of 2.3767, 0.79239 and 1.8769 eV for Fe, Ni and Co respectively. The site moments may be compared with the total spin per unit cell, which is 1.10, 0.314 and 0.815 respectively. The difference between the local and unit cell spin may be regarded as an inherent uncertainty in the method, which crucially relies on mapping an itinerant system to a localized spin model. The calculated magnon dispersions exhibit a significant effect of the renormalization in Ni and Co due to the rather small exchange splitting in these compounds, whereas the renormalization is less pronounced in Fe. The renormalized MFT dispersions are in excellent agreement with magnonenergies obtained from time-dependent density functional theory (TDDFT) [39], whereas the bare MFT expressions underestimate the TDDFT results. The actual magnetic response of these systems is prone to rather strong Landau damping—in particular when approaching the Brillouin zone boundary—and the definition of a magnon energy from the TDDFT spectral function becomes ambiguous in some parts of the Brillouin zone [39]. The MFT approach will, however, by construction yield a magnon energy at any $\mathbf{q}$ and the noisy features of the magnon dispersions of Fig. 2 are related to nontrivial (not just Lorentzian broadening) interactions with the Stoner continuum. Nevertheless, we may use the MFT magnon energies to evaluate critical temperatures based on the different decoupling schemes discussed above, keeping in mind that the Heisenberg modelling of such systems is rather severely limited by the itinerant nature of the compounds.